1) For the chemical reaction, 2CH4(g) <-->C2H2(g) + 3H2(g) , K= 0.165 at1752o C. What is Kp for the reaction at thistemperature?
2) For the reaction, N2(g) + 3H2(g)<--> 2NH3(g) , Kp = 4.95 x10-3 at 270oC. What is K for the reaction atthis temperature?
Answer #1
1) For the chemical reaction, 2CH4(g) <-->C2H2(g) + 3H2(g) , K= 0.165 at1752o C. What is Kp for the reaction at thistemperature?
We know that Kp = Kc * (RT) Δn
Where T = Temperature = 1752 o = 1752 + 273 = 2025K
R = gas constant = 0.0821 L atm / mol - K
Δn = No . of moles of gaseous products - No . of moles ofgaseous reactants
     = (1+3) - 2
     = 2
Plug the values we get Kp = 0.165 * ( 0.0821 * 2025 )2
                                        = 4.5605*10 3
2) For the reaction, N2(g) + 3H2(g)<--> 2NH3(g) , Kp = 4.95 x10-3 at 270oC. What is K for the reaction atthis temperature?
We know that Kp = Kc * (RT) Δn
Where T = Temperature = 270 o C = 270 + 273 =543K
R = gas constant = 0.0821 L atm / mol - K
Δn = No . of moles of gaseous products - No . of moles ofgaseous reactants
     = 2 - ( 1+3)
     =-2
Plug the values we get K = Kp *  (RT)-Δn
                                      = 4.95 x 10-3 * ( 0.0821 *543 ) 2
                                        = 9.837